OpenStudy (anonymous):

Simplify the expression and eliminate any negative exponent(s). Assume that all the letters denote positive numbers.

OpenStudy (anonymous):

\[(\frac{ x^{-2/3}}{ y ^{1/2} })(\frac{ x^{-2}}{ y ^{-3} })^{1/6}\]

OpenStudy (anonymous):

this is what i did \[(\frac{ x ^{-2/3} }{ y ^{1/2} })(\frac{ x ^{-3} }{ y ^{-2} })\]

OpenStudy (anonymous):

and please show your work

OpenStudy (anonymous):

@austinL

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

exponents raised to other exponents = multiplication. I assume you have the basic rules. lemme work through this

OpenStudy (anonymous):

ok thanks buddy, i completely have no idea of what the next step should be

OpenStudy (anonymous):

maybe i need -2/3 exponent to become positive and that is by multiplying it with the other expression...instead of dividing..

OpenStudy (bibby):

add the exponents

OpenStudy (anonymous):